Two numbers are in the ratio 5:8. If their difference is 36, what is the larger number?
Answer by ikleyn(52786)

.

The numbers are 5x and 8x,  and

    8x - 5x = 36,


giving

      3x   = 36

       x   = 36/3 = 12.


The numbers are 5*12 = 60 (the smaller) and 8*12 = 96 (the larger).


ANSWER.  The larger number is 96.

Solved.